OpenStudy (anonymous):

Find the product: (x + 4)4

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf (a+3)3=3a+9 }\)

OpenStudy (solomonzelman):

unless you meant "4" as an exponent.

OpenStudy (anonymous):

If it's an exponent, (x+4)^4 = (x+4)(x+4)(x+4)(x+4)